### v2

# Deconstructing The Ethics of Large Language Models from Long-standing Issues to New-emerging Dilemmas: A Survey

Abstract
Large Language Models (LLMs) have achieved unparalleled success across diverse language modeling tasks in recent years. However, this progress has also intensified ethical concerns, impacting the deployment of LLMs in everyday contexts. This paper provides a comprehensive survey of ethical challenges associated with LLMs, from longstanding issues such as copyright infringement, systematic bias, and data privacy, to emerging problems like truthfulness and social norms. We critically analyze existing research aimed at understanding, examining, and mitigating these ethical risks. Our survey underscores integrating ethical standards and societal values into the development of LLMs, thereby guiding the development of responsible and ethically aligned language models.
keywords:
Large Language Models, Trustworthy, Ethics, Survey
1 Introduction
$\dagger$ All authors contributed equally.
In the past few years, the field of artificial intelligence (AI) has witnessed a surge in the development of large language models (LLMs). These advanced computational language models have demonstrated remarkable performance across a spectrum of language modeling tasks [46, 259, 292, 344, 354, 353, 192, 312]. Their capabilities are exemplified in natural language generation [38, 47, 209], where they can create coherent and contextually relevant text, question answering [15, 337, 357], where they effectively retrieve or infer information in response to queries, and complex reasoning tasks [119, 131, 334, 310], which involve navigating through intricate problem-solving processes. Despite these advancements, LLMs have also raised substantial ethical concerns. As these models become increasingly integrated into daily life, addressing these ethical challenges becomes paramount. The concerns are multifaceted, encompassing issues such as privacy [307], copyright, robustness [335], bias, and the potential for misuse. Given their ability to understand and generate human-like responses, there’s a growing discourse on ensuring these responses are not only accurate but also ethically aligned with societal norms and values.
In response to ethical concerns, substantial research is focusing on the ethical implications of LLMs. Scholars aim to identify, examine, and mitigate potential risks, guiding the development of more responsible AI systems [52]. This effort ensures LLMs are designed and deployed to maximize benefits and minimize harm, serving the public good ethically and effectively. The realization of these objectives hinges heavily on access to large-scale high-quality corpus and textual datasets. However, collecting the data may bring ethical issues, such as privacy, copyright, and bias [307]. These ethical issues are long-existing and still challenging. Besides, some new ethical issues emerge as LLMs develop. For example, there is a growing concern over the potential for LLMs to produce inappropriate responses to unethical queries. To avoid this issue, alignment techniques are developed to align the answers with human values [178]. Similarly, the phenomenon of model-generated content that lacks factual basis, often referred to as “hallucinations”, presents another ethical concern [339]. Furthermore, some new issues may emerge during the applications of LLMs, such as law and regulatory compliance [149]. To illustrate, we outline the significant ethical issues for each subsection as follows:
  • Privacy issues brought by LLMs include but are not limited to memorization (or data leaking), and privacy attacks. To provide a comprehensive review of ethics issues in privacy concerns, we first introduce existing privacy issues and their challenges and further provide two aspects of alleviating methods, differentiable privacy LLMs and emerging methods of preserving privacy.
  • Copyright concerns may be raised in LLM-generated content. We chronologically introduce two main technology arms of copyright - backdoor and watermark - to demonstrate their expansion and diffusion. For example, our introduction ranges from protecting web texts by HTML coding to preserving general texts on embodied watermarks, and from protecting the outputs to safeguarding the generative model and datasets, etc.
  • Fairness problems, such as societal biases in the training data of LLMs, may cause harm to marginalized communities, like prejudices, stereotypes, and discriminatory attitudes. To provide inclusive and equitable LLM-based services, it is critical to prevent LLMs from unintentionally perpetuating or amplifying these biases when generating responses.
  • Truthfulness of LLMs may be undermined by hallucination and sycophancy issues. Specifically, hallucination problems may inadvertently result in generating false information that appears credible, whereas sycophancy issues may amplify human preference rather than correct response. Addressing these two concerns is crucial to maintaining the trust and credibility of LLM technologies.
  • Social Norm plays a pivotal role in our society. However, LLMs may produce toxic content due to the contamination of train data. Alignment is one of the crucial techniques to address toxicity. In this survey, we discuss the motivation, characteristics, and recent advancements in alignment techniques, which are critical in the development and deployment of LLMs.
  • Law and Regulatory Compliance for LLMs are essential in our society as worldwide governments urgently promote AI-related legislation, such as the EU AI Act, to ensure that the utilization of AI tools aligns with ethical standards.
Figure 1: Main category in this survey paper.
In this survey, we aim to investigate ethical issues in the development of LLMs and propose a new taxonomy to help readers better understand the ethical issues and corresponding techniques that are proposed to solve these issues. Specifically, we categorize the ethical issues as longstanding problems and new-emerging problems. In the former category, we mainly discuss the ethical problems in 1) data privacy, 2) copyright, and 3) fairness. For the latter category, we are interested in the topic of truthfulness and social norms. Also, We introduce the law and regulatory compliance in the era of LLMs. To better illustrate our proposed taxonomy, we present the overall hierarchy in Figure 1. In brief, we summarize our contributions in this survey as follows:
  • We systematically summarize and categorize existing ethical issues into two main categories: 1) we discuss longstanding problems of data privacy, copyright, and fairness; 2) we investigate new-emerging problems that are pertinent to LLMs, including truthfulness and social norms, and further discuss the design and requirement of law and regulatory compliance in guiding future explorations.
  • We introduce the existing issues and mitigation strategies, and further present the hierarchy for each category in Figure 2 and Figure 5.
  • We discuss the future research directions for each section of the ethical issues.
The subsequent sections of this paper are structured as follows: Section {2} delves into enduring ethical dilemmas predating the advent of LLMs, while Section {3} introduces newly emergent ethical concerns in the era of LLMs.
2 Persistent Ethical Issues
In this section, we present the longstanding ethical problems predating the advent of LLMs. These include 1) data privacy, 2) copyright, and 3) fairness. The hierarchy is displayed in Figure 2.
Figure 2: The hierarchy of longstanding ethical problems in Section 2. We list corresponding mitigation strategies for each sub-category.
2.1 Data Privacy
2.1.1 Privacy: Issues and Challenges
Data privacy has long been a concern, but there is a growing consensus that while Large Language Models (LLMs) offer impressive capabilities, they also raise significant data privacy issues today. In this section, we first introduce issues and potential challenges and then discuss major solutions regarding these issues (e.g. Section 2.1.2 deferentially private LLMs and other emerging techniques in Section 2.1.3). The concerns in privacy could be mainly summarized in twofold, memorization and privacy attacks as illustrated in Figure 3.
Figure 3: Data privacy issues & challenges detailed categories and mitigation methods.
Memorization. All machine-learning (ML) models, including LLMs, inherently memorize to some extent, as they learn by observing and recalling training data. However, this problem becomes severe when it comes to LLMs because of its tremendous size and capacity. We list the main aspects of risk factors that may affect the memorization issues.
  • Model capacity: The capacity of a model significantly impacts its memorization ability. Larger models, as shown by [44] and [268], tend to memorize more data and do so at a faster rate. This memorization is not directly linked to model performance, as shown by comparing GPT-2 and GPT-Neo models. The trend suggests that neural networks’ capacity to memorize is substantial and growing, outpacing the size increase of language datasets.
  • Dataset scale: Research on dataset size and memorization reveals contrasting findings. Li et al. [156] discovered that larger datasets lead to less memorization, evidenced by a decline in canary extraction success over training time. Conversely, Biderman et al. [30] found that points memorized early in training tend to be retained in fully trained models, suggesting persistent memorization despite dataset size.
  • Sample duplication is a key factor in memorization for Large Language Models (LLMs). Lee et al. [150] observed that data duplication in large web datasets follows a power law, with a small fraction of data being highly duplicated. This duplication significantly increases memorization, as models trained on deduplicated datasets exhibit much lower rates of outputting memorized text. Kandpal et al. [138] further demonstrated that sequences repeated in the dataset are generated far more frequently by LLMs. Despite this, memorization still occurs even with little or no data duplication, indicating other contributing factors to memorization beyond mere duplication.
  • Prompting significantly affects memorization in Large Language Models (LLMs). Mccoy et al. [196] observed that longer generated sequences tend to produce more novel content, reducing memorization. Conversely, longer prompts increase memorization for a constant n, as shown by [44]. Additionally, specific token types, like nouns and numbers, are memorized faster than others, such as verbs and adjectives. Kharitonov et al. found that larger subword vocabularies in tokenizers lead to increased memorization, possibly due to reduced sequence length making it easier for models to memorize [141].
Privacy Attack. The robustness of Large Language Models (LLMs) may be weakened by privacy attacks. We list three scenarios that may bring privacy risks to the robustness issues of LLMs as follows.
  • Membership inference attacks (MIAs) have been recently studied on language models (LMs). While LMs are generally resistant to simple probing, they are vulnerable to sophisticated MIAs. Threshold attacks on embedding models by [256] and perplexity-based attacks on GPT-2 by [43] revealed privacy risks. Reference model-based attacks like [206] improved detection accuracy, while Mattern et al. developed a neighbor comparison framework without database access [193]. Additionally, Tople et al. exploited model updates for data exposure [272], and some works used various methods for successful MIAs [202, 111, 200]. Shadow model attacks also proved effective, with research by [2, 42] showcasing risks even in pre-trained datasets. These findings highlight the evolving nature and potential privacy concerns of MIAs in LMs.
  • Training data extraction is a privacy attack enabling adversaries to retrieve sensitive data using query access. Carlini et al. pioneered this method, involving generating candidate targets, applying a membership inference attack (MIA), and selecting top-k candidates [43]. Their experiments on GPT-2 demonstrated the feasibility of extracting training data, including sensitive personal information. Subsequent research by [326, 341] introduced improvements in candidate generation and MIA processes, significantly enhancing extraction precision. Nasr et al. extended these attacks to production LMs like ChatGPT and open-source models, revealing higher memorization levels than previously understood [214]. This line of research underscores the potential privacy risks inherent in LMs and the effectiveness of training data extraction attacks.
  • Attribute inference attacks represent a privacy risk for LLMs, though less researched than membership inference and training data extraction attacks. Staab et al. conducted a comprehensive study of this risk by using LLMs to infer personal attributes from public user data like online forum posts [258]. They tested various LLMs, including GPT-4, and used a database of annotated Reddit profiles to assess accuracy in predicting attributes like age, education, and income. GPT-4 achieved a high accuracy rate of 84.6% across all attributes. This study highlights that while attribute inference attacks are a potential privacy risk with LLMs, such risks are not exclusive to these models but could be exacerbated by their efficiency.
2.1.2 Differentially Private LLMs
Differential privacy (DP) [72] emerges as the primary scheme to address data privacy concerns. Acknowledged as de facto golden standard, differential privacy provides mathematical rigor to the algorithms involving sensitive information to be protected. Essentially, an algorithm is differentially private if the output distribution is relatively close, tailored by certain privacy parameters whether an individual’s data is present or not in the dataset. More formally, we denote differential privacy as follows.
Definition 1
(Differential Privacy) Given two databases $Y$ and $Y^{\prime}$ that are identical except for one data entry, a randomized algorithm $\mathcal{M}$ is $(\epsilon,\delta)$ differentially private if for any measurable set $A$ in the range of $\mathcal{M}$, $\Pr\left[\mathcal{M}(Y)\in A\right]\leq e^{\epsilon}\Pr\left[\mathcal{M}(Y^{\prime})\in A\right]+\delta$.
An ideal DP algorithm protects the data privacy with the given $(\epsilon,\delta)$ guarantee meanwhile minimizing the performance degradation compared to the ground truth. In the realm of machine learning, the mainstream technique of applying DP is Differentially Private Stochastic Gradient Descent (DP-SGD) [1], where the gradient is first clipped and then perturbed with Gaussian noise at each step of the optimization. Most existing DP techniques for language models are developed upon DP-SGD. Before delving into details, one caveat is DP requires a primitive definition on the ‘resolution’ of privacy preservation, that is, where does one data entry (Definition 1) zoom into? For NLP tasks, one data entry could be data of one user (resp. user-level), a sentence (resp. sequence-level), or a word/token (resp. token-level), etc. In many cases, user-level DP is captured by local DP while the rest falls in centralized DP approaches. Apparently, various scopes of the DP concept are impactful on algorithm design and performance evaluation. We therefore include this front for each work if the context is clear.
In the pre-LLM era, techniques involving differential privacy can be categorized into DP (pre)training and DP fine-tuning. As language models scale up, training and fine-tuning with large loads can be prohibitively expensive in certain scenarios. DP inference, as a new paradigm, harmonizes with new techniques in LLMs such as in-context learning and prompt tuning, etc. Therefore we focus on DP inference as the main remedy of the data privacy issue in the LLM era.
Pre-LLM Era. We first explore existing methods that employ DP training, where a language model is usually trained from scratch using variants of DP-SGD. An early attempt, DP-FedAve [197] dates back to the ante-transformer era. It targets recurrent language models and introduces a DP optimization technique inspired by a federated averaging algorithm. Consequently, differential privacy is defined on the user level. To improve the privacy-utility trade-off, a later work, Selective DP-SGD [250] introduces the concept of selective differential privacy, which provides focused protection for sensitive attributes only in one training example. Note that this method only applies to RNN-based language models. Moving forward to pre-trained transformer language models, two closely related works [112, 10] improve DP-SGD and train BERT with DP guarantees. Both consider the protection level as item-level, which is one training example containing several words. The latter work [10] focuses on training heuristics that bring more efficiency and can be implemented on BERT-large.
Fine-tuning language models for downstream tasks also provokes privacy issues on domain-specific data. Even though differential privacy (DP) techniques for model fine-tuning emerged before the advent of large language models (LLMs), they continue to hold potential in the LLM era. Historically, these techniques have been tested primarily on models with million-scale parameters. Recent advancements in DP fine-tuning [324, 182, 158] suggest that larger models might offer improved trade-offs between privacy and utility for such tasks, as highlighted in concurrent studies Further, Yu et al. [324] developed an innovative optimization approach for example-level DP that eliminates the need for generating per-example gradients in DP-stochastic gradient descent (SGD), thereby conserving memory. In a similar vein, Li et al. [158] consider user-level DP and claim that parameter-efficient fine-tuning can achieve impressive efficiency while keeping good utility. Experiments are carried out on RoBERTa families [179] and GPT families [234, 235, 38]. With a similar aim for efficiency, DP-decoding [191] proposes a simple perturbation mechanism applied to the output probability distributions, which is sufficient for privacy guarantee due to the post-processing lemma [72].
LLM Era. LLMs demonstrate compelling capabilities such as in-context learning merely within the inference stage. Privacy-preserving approaches lying in this category bypass the projection of DP-SGD and commonly add perturbation to more accessible information sources such as prompts or embeddings, leaving LLMs parameters frozen. With respect to in-context learning, two works [304, 265] emerge with a similar scheme of ‘divide-and-privately-aggregate’, however, considering different privacy levels. DP-ICL [304] aggregates the LLM responses for each group of exemplars with differential privacy. Two mechanisms are proposed for private aggregation: embedding space aggregation and keyword space aggregation. DP-ICL is on the user level while the later work [265] zooms into the example level, the aggregation algorithms are based on the Gaussian mechanism and exponential mechanism and applied to exemplars in sensitive datasets. Another work on privacy-preserving prompt tuning called RAPT [159] also privatizes source datasets with DP guarantees, where tokens are reconstructed with randomized mechanisms, and then trained jointly with the downstream tasks. Last, we include three recent methods that apply DP by adding perturbation to embeddings. DP-forward [70] devises an analytic matrix Gaussian mechanism that perturbs the embedding matrices in the forward pass of language models. Split-N-Denoise [189] further provides a framework where the embeddings are first perturbed on the user side and then transmitted to the server. A denoising module can be trained to produce outputs given noisy responses from the server LLMs. Both works consider local DP. Shortly after, InferDPT [270] moves to document-level DP that protects sensitive information in prompts for black-box LLM inference. The proposed pipeline contains a perturbation module based on an exponential mechanism and an extraction module that selects coherent and consistent text from the perturbed generation result.
2.1.3 Other emerging methods
There also exists a diverse array of alternative methods that primarily focus on two key areas: privacy preservation within distributed frameworks and the processing of data in ways that safeguard sensitive information. Distributed frameworks, such as federated learning, offer a decentralized approach where data processing and model training occur locally on user devices, thus minimizing the exposure of sensitive data [135, 331]. This approach contrasts with differential privacy, which typically adds noise to datasets or queries to prevent the identification of individual data points. Federated learning addresses privacy concerns by ensuring that sensitive data remains on the user’s device. Only the model updates, which are less likely to contain personally identifiable information, are shared. Several federated learning algorithms have been proposed for LLM training [309], fine-tuning [115, 342, 146, 93], and few-shot learning [130]. However, federated learning can still be vulnerable to adversary attacks that target private text  [93, 19, 75, 56, 239]. Future efforts could aim to defend by leveraging training strategies such as fine-tuning on private datasets [93].
Furthermore, advanced data processing techniques, including secure multi-party computation (SMPC) [87, 60], enable the manipulation of encrypted data without revealing its contents. These methods provide robust privacy guarantees and are particularly advantageous in scenarios where data cannot be shared openly due to privacy concerns or regulatory constraints. SMPC provides higher privacy guarantees than federated learning methods as the latter exposed the shared model parameters across participating parties which could potentially expose information about the data [211, 277, 332, 73]. As a trade-off, SMPC may face challenges that could impact the efficiency and effectiveness of the model. The computational complexity of SMPC, due to its cryptographic operations, often results in slower processing times and increased resource consumption, particularly for LLMs. Therefore, existing approaches aim to speed up SMPC inference for common network architectures such as transformers in LLMs [152, 92, 346, 69, 116, 99, 49] or adapting existing model frameworks to enhance efficiency  [330, 166]. For a deeper dive into SMPC defense strategies for LLMs, we direct the readers to [159].
Furthermore, machine unlearning and data sanitization have just started to gain attention, each addressing privacy concerns at different stages of data handling. Machine unlearning is a process designed to efficiently and effectively remove specific data from an already trained model. This is particularly relevant in scenarios where users wish to retract their data due to privacy concerns or in compliance with regulations like General Data Protection Regulation (GDPR) [280], which includes the ‘right to be forgotten’. For large language models, this involves retraining or adjusting the model in a way that the influence of the specific user’s data is negated, without having to retrain the model from scratch [316, 226, 300]. Data sanitization refers to modifying data to remove or alter sensitive information before being used for training models [138, 123]. However, a major limitation is the potential for excessively removing training data [31], which can be a future research focus.
2.2 Copyright
Copyright has been a long-existing legal issue in the natural language domain [23] that calls for research on encoding imperceptible and indelible signatures on plain texts to protect the property of authorship [8]. In literature, as an information hiding application [22], the traditional techniques extend from steganography [59] to watermarking [253]. In the language model era, copyrights preserving techniques further develop to protect the model rather than sorely the data, where backdoor [50, 89, 62, 160, 81, 188] and watermark [144] are two main streams. The hierarchy in this section is portrayed in Figure 4.
Figure 4: Copyright methods.
2.2.1 Backdoor
In backdoor attacks [89, 216, 137, 79, 16, 186, 306, 185, 187], the attacker constructs poisoned samples by adding an attacker-defined trigger to a fraction of the training samples and changing the associated labels to a specific target class. A backdoor can be injected by training the model with a mixture of clean and poisoned samples. A backdoor-compromised model functions normally with clean inputs but exhibits abnormal behavior when presented with inputs containing a specific trigger. By embedding a unique trigger pattern within a model through a backdoor, a distinct relationship between the trigger and the target label is established. In classification tasks, the presence of the trigger will consistently induce the model to predict the corresponding target label. These properties can be used to signify the model’s ownership or origin, particularly in situations where the model is not accessible, such as in a black-box setting.
Pre-LLM Era. Adi et al. first introduce that Backdoor can be used as watermarks for ownership verification [6]. To avoid detection, Xiang et al. propose a semantic and robust watermarking scheme for natural language generation (NLG) models that utilize unharmful phrase pairs as watermarks for intellectual property (IP) protection [305]. He et al. use lexical replacements of specific words to demonstrate ownership for LLMs deployed through APIs [107]. In addition, large pre-trained language models (PLMs) require fine-tuning on downstream datasets, which makes it hard to claim the ownership of PLMs. Gu et al. show that PLMs can be watermarked with a multi-task learning framework by embedding backdoors, making watermarks difficult to remove even after fine-tuning the models on multiple tasks [88]. Shokri et al. investigate membership inference attacks on machine learning models trained by commercial ”machine learning as a service” providers such as Google and Amazon, determining if a data record was part of the training dataset. [252]. Liu et al. present a novel watermarking technique using a backdoor-based membership inference approach via marking a small subset of samples for data copyright protection in the black-box setting [180].
LLM Era. Copyright protection of LLMs has become crucial due to the substantial training cost associated with these models. Liu et al. indicate that LLMs are vulnerable to model extraction attacks, wherein attackers can copy the model using query texts and returned embeddings, potentially building their own LLMs and causing significant losses for the original model owners [181]. EmbMarker [228] proposes to implant backdoors on embeddings of LLMs. Specifically, it selects moderately frequent words as triggers, defines a target embedding as the watermark, and uses a backdoor function to embed it. Lucas et al. propose an attack to identify trigger words or phrases by analyzing open-ended generations from LLMs with backdoor watermarks [184]. It is shown that triggers based on random common words are easier to detect than those based on rare tokens.
Discussion. We suggest that the exploration of stricter settings is necessary. For example, in most research, data owners have access to the percentage of their data within the total training set, which necessitates knowledge of tasks associated with PLMs. Hence, how to adapt the backdoor-based methods for stricter settings in copyright protection remains an open direction. In addition, as the field of backdoor-based copyright protection advances, an increasing number of tailored model-stealing techniques are being studied, such as knowledge distillation [110]. It is essential to explore the resilience of backdoor-based algorithms against potential attacks that adversaries may employ. Finally, the effectiveness of backdoor-based copyright protection for LMs still lacks a comprehensive theoretical framework. The clarity of such a framework remains an open question in this field.
2.2.2 Watermark
Watermarking aims to conceal invisible signatures in plain text and be extractable for future examination, which has been a solution to copyright protection for a long time. However, due to the discrete nature of natural language, the capacity, robustness, and invisibility are more challenging to achieve than other media like images, audio, and videos. Brassil et al. first comprehensively introduce mechanisms for marking and decoding watermarks specifically for the texts to prevent illegal copies [36]. In the past two decades, digital watermarking on format, scanned image, frequency of words, syntactic, and semantics has been proposed [8]. The trend of watermarking renews in the era of LLMs for detection to prevent abuse [144]. The possibility of adding human-imperceptible signatures during the decoding stage of LLM is under wide exploration.
Pre-LLM Era. Watermark is first concerned as an information hiding technology for a small amount of information [231]. Mir et al. apply this technique to protect the copyright of web content [205]. Early approaches of watermarking include text-meaning representations of sentences for information hiding by syntactic rules [12], watermarking on the format of documents by vertical and horizontal line-shifting [37], watermarking by inserting zero-width control characters in Hyper Text Markup Language (HTML) [9], watermarking on semantics by synonyms substitution [24, 271], and zero-watermarks by using word length [127] and contents of text [128].
LLM Era. Watermarking at the current stage focuses more on the model schemes for watermarked generation. As pioneers, Kirchenbauer et al. propose an LLM watermarking algorithm by adding token-level bias in the decoding stage [144]. Kuditipudi et al. design a distortion-free watermark to preserve the original distribution of LLM during watermarking [147]. Ren et al. consider the semantic embedding in hashing tokens [245] and Fu et al. concern semantic word similarity to enhance the robustness [76]. Yoo et al. embed multi-bit information into the watermark, which succeeds traditional steganography [322]. They inject the watermark via word replacement after initial generation, which is further integrated into one stage by [288]. Christ et al. propose a computationally undetectable watermark theoretically if the secret key is inaccessible [54]. Liu et al. propose a private watermark utilizing separated neural networks respectively for generation and detection [167]. The aforementioned works focus more on the token level, while there are emerging works focusing on a higher-level perspective. Hou et al. introduce a sentence-level semantic watermark that aims at periphrastic robustness [113, 114]. For applications, some works mention the importance of watermarking the ownership of datasets via inference [190, 172]. Yao et al. introduce copyright protection for prompts via watermarking [315].
Discussion. One of the main challenges for watermarking is its popularization and the opening of corresponding detection methods and configurations. Hopefully, this requires administrative oversight from government and industry associations. US Federal, China, and Europe have mentioned potential proposals in some of the government documents, e.g., Interim Measures for Generative Artificial Intelligence Service Management of China, Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence of the US, and the European Union’s AI Act. Moreover, the definition and notion of authorship are slightly ambiguous as human-LLM collaboration and multi-agent generation are becoming mainstream. Tripto et al. discover that literate studies have contrasting perspectives on whether authorship remains the same after paraphrasing, as paraphrasing deviates the style of text dramatically [276]. Meanwhile, further improvement on the watermark’s robustness to attack [290], generalization to short contents, reduction of impact on text quality, and differentiation to direct machine-generated text detection [83, 207, 177, 175] are worth exploring.
2.3 Fairness
LLMs inherit and potentially amplify societal biases present in their training data, which can perpetuate harm against marginalized communities [21]. Fairness issues can be in various NLP tasks, such as text generation [164, 314], machine translation [199], information retrieval [243], natural language inference [65], classification [210, 358] and question-answering[66, 224]. They can be influenced at different stages of the LLM deployment cycle, including training data, model architecture, evaluation, and deployment, which has been thoroughly explored by [201, 264]. Fairness and bias definitions are crucial for understanding the challenges and addressing them in LLM, as they provide a foundation for developing and evaluating mitigation strategies.
We consider the following fairness definitions. Group Fairness focuses on disparities between social groups, which is defined as requiring parity across all social groups in terms of a statistical outcome measure [53, 100, 170, 136, 102, 319, 333]. Individual Fairness is defined as the requirement that individuals who are similar in a task should be treated similarly [71, 105]. It involves a measure of similarity between distributions of outcomes [104, 106]. Social Bias is defined as encompassing disparate treatment or outcomes between social groups arising from historical and structural power asymmetries [20, 32, 61]. In NLP, this includes representational harms (like misrepresentation [254], stereotyping [4], disparate system performance [33, 356], derogatory language [29], and exclusionary norms [21]) and allocational harms (such as direct and indirect discrimination [74]). In the following subsections, we study this crucial issue by categorizing, summarizing, and discussing research on measuring and mitigating social bias in LLMs.
2.3.1 Mitigation Strategy
Bias mitigation in traditional machine learning involves pre-processing data to reduce bias, altering algorithms during training (in-processing), and adjusting outputs post-training (post-processing). In the LLM era, similar strategies are employed: pre-processing techniques reduce bias in training data and prompts, in-training methods modify training procedures and model architecture, intra-processing approaches generate debiased predictions during inference, and post-processing techniques address bias in outputs, particularly for black-box models.
Pre-LLM Era. As machine learning models are increasingly deployed in critical domains [145, 311, 352, 118, 351], addressing bias to achieve fairness has become essential. Traditional bias mitigation approaches are categorized into three main strategies. Pre-processing techniques aim to modify the data by reducing inherent biases [63]. For example, Pessach et al. [229] suggest a pre-processing mechanism to enhance fairness in private collaborative machine learning scenarios [340, 48]. In-processing methods involve altering learning algorithms to eliminate bias during model training [308]. Berk et al. [25] introduced fairness regularizers for linear and logistic regression models to ensure both group and individual fairness. Post-processing techniques are applied after training, adjusting model outputs to enhance fairness [143]. Petersen et al. [230] developed a general post-processing algorithm that ensures individual fairness by utilizing graph Laplacian regularization [297], framing the challenge as a graph smoothing problem.
LLM Era. Bias mitigation techniques in LLMs also follow a similar pattern and can be categorized into four groups based on their application at different stages of the LLM workflow: pre-processing, in-training, intra-processing, and post-processing [78].
Pre-processing Mitigation. These techniques aim to reduce bias in training data and prompts before training. There are various methods in this category. The first method involves neutralizing bias by adding new examples to extend the representation of underrepresented social groups. Techniques include counterfactual data augmentation [232, 85], selective training example substitution [194, 328], etc. The second method applies instance weighting to balance class influence to increase the impact of existing biased examples [98, 220], and applies reweighting token probabilities in pre-trained models during knowledge distillation to prevent bias transfer [64, 325]. The third method focuses on creating new examples adhering to specific characteristics, like collecting high-quality examples to steer the model towards desired output [260, 142], and generating word lists associated with social groups [94]. The fourth method performs instruction tuning by adding textual instructions [213], static tokens [183], or trained prefixes [157, 176] to reduce bias in the data. There is also one line of work involving altering contextualized embeddings to remove bias [240, 124].
In-training Mitigation. These mitigation techniques focus on modifying the training procedure to reduce bias. The first method of this category focuses on altering the model’s structure (i.e., integrating debiasing adapter modules [117]), and using demographic-specific encoder [98]. The second method focuses on disrupting the association between social groups and stereotypical words. This is typically achieved by modifying the loss function applied on various model layers like the embedding layer [173, 223], attention layers [77, 13], and token generation stage [233, 109]. Additionally, new training paradigms are proposed, such as contrastive learning [219, 161], adversarial learning [97, 242], and reinforcement learning [174, 18]. The last method focuses on efficient fine-tuning procedures that freeze most pre-trained model parameters, and only update those potentially related to bias [323, 291, 285, 293].
Intra-processing Mitigation. These approaches modify a trained model’s behavior without additional training to generate debiased predictions during inference. There are mainly four types of methods. The first method adds restrictions during token search decoding to prevent biased outputs [249, 198]. The second method adjusts token distributions to enhance output diversity or sample less biased outputs [58, 96]. The third method redistributes the model’s attention to less stereotypical aspects [327]. The last method implements standalone networks with original models for specific debiasing tasks, such as reducing gender or racial biases [101].
Post-processing Mitigation. The techniques address bias in generated outputs, especially relevant for black-box models with inaccessible training data or internal processes. The techniques can be mainly categorized into two types. The first type of method uses explainable machine learning to identify biased tokens and replace them with unbiased alternatives [269, 67], or employing protected attribute classifiers for this purpose [108]. The second type of method treats the mitigation as a machine translation task, transforming biased sentences into unbiased ones [126, 262, 278].
2.3.2 Measurements on Fairness
Measurements on LLMs’ fairness are generally categorized into three types, based on the model elements they analyze: embeddings, probabilities, and generated texts [78].
Embedding-based Metrics involve calculating the distances within the embedding space between neutral terms, like job titles, and identity-specific terms, such as gender pronouns [40, 195, 91, 68]. In an unbiased model, the distance between neutral and diverse social group terms should be comparably similar in the embedding space.
Probability-based Metrics involves prompting the model with template sentences that have variations in their social group terms. The main focus is on comparing the probability distribution of predicted tokens, conditioned on the rest of the input [295, 7, 212, 139, 103]. A model that demonstrates no bias should yield consistent probability distributions for attributes, regardless of any alterations in the protected characteristics.
Generated Text-based Metrics evaluate the text produced by LLMs and are particularly valuable for models treated as ’black boxes’, where direct access to probabilities or embeddings is not feasible. This category includes three distinct types of metrics: Distribution Metrics assess the frequency distribution of tokens related to various social groups in the generated text [237, 35]. Classifier Metrics employ an auxiliary model to estimate the degree of social bias present in the text produced by the LLM [121, 255]. Lexicon metrics involves comparing each word in the LLM’s output against a pre-established list of terms to calculate a biased score [218, 66]. An unbiased and fair model should output similar distributions, or biased scores for different social groups or neutral terms.
Discussion. To effectively mitigate bias in LLMs, it is essential to adopt a comprehensive approach that leverages the strengths of various bias mitigation strategies. Specifically, pre-processing techniques should be employed to neutralize biases at the source, ensuring that the data used to train the LLM is as unbiased as possible. Subsequently, in-training mitigation strategies can be implemented to further refine the training process of the LLM, improving its ability to produce fair and unbiased outputs. Finally, during the model’s deployment phase, both intra-processing and post-processing measures could be applied to minimize the risk of generating biased content. By combining these methods, we can create a robust framework that significantly reduces the likelihood of bias in outputs, fostering a more equitable and fair use of LLMs.
3 New-emerging Ethical Issues
In this section, we introduce the new-emerging ethical problems related to truthfulness and social norms that emerged during the era of LLMs. We also discuss the progress of regulatory compliance as the development of LLMs. The hierarchy in this section is portrayed in Figure 5.
Figure 5: The hierarchy of new-emerging ethical problems in Section 3. We list the ethical issues and corresponding mitigation strategies for each sub-category.
3.1 Truthfulness
Truthfulness in LLM is a critical concern due to issues like hallucination and sycophancy, both of which compromise the reliability and ethical deployment of these technologies. Hallucination refers to the generation of factually incorrect or misleading information, which can severely compromise the reliability of LLMs in critical applications such as medical diagnosis or legal advice. Sycophancy, on the other hand, manifests as an undue eagerness to affirm user opinions, potentially leading to biased or overly positive responses that may not reflect accurate information. In extreme scenarios, such biased models may not only reinforce users’ pre-existing beliefs but may also promote actions that are ethically or legally questionable.
Addressing these issues is crucial for the integrity and utility of LLMs. Developing mechanisms to ensure that LLMs consistently maintain factual accuracy and neutrality is essential, especially for their integration into decision-making processes where trust and objectivity are paramount.
3.1.1 Hallucination
Large language models tend to produce hallucinations where the models generate contents that deviate from the input, contradict existing contexts, or misalign with universally accepted world knowledge [155, 287, 320, 84]. An example is shown in Figure 6. Such phenomena pose significant challenges, particularly when considering the reliability and trustworthiness of LLMs in critical applications. We delve into the underlying causes, manifestations, and potential mitigation strategies for hallucinations in LLMs.
Figure 6: An example of hallucination. The LLM correctly identifies Thomas Edison as a key figure in the invention of the lightbulb in 1879. However, the model also fabricates information about Benjamin Franklin inventing an electric light in 1802, which is inaccurate. Benjamin Franklin is well-known for his experiments with electricity, particularly the kite experiment, but he did not contribute to the invention of the lightbulb. The model ”hallucinated” this fact, likely by confusing Franklin’s work with electricity with the development of the lightbulb.
Underlying Causes. The primary causes of hallucinations in LLMs can be broadly categorized into data quality, model architecture, and algorithmic limitations:
  • Data quality: Models trained on datasets with inaccuracies, biases, or limited scope are more susceptible to hallucinations. Such data compromises the model’s representation of reality, leading to outputs that significantly deviate from correct input, contradict established contexts, or misalign with universally acknowledged facts.
  • Model architecture: Despite their complexity, current LLMs lack true comprehension similar to human understanding. They rely on patterns in datasets rather than in-depth content understanding for response generation, which can produce structurally coherent but content-flawed outputs [14, 168, 133]. The size of models also poses risks. While it enables learning from diverse data, it also increases the likelihood of incorporating flawed information [289, 171, 321]. Overconfidence in outputs caused by insufficient human oversight, sparse alignment examples, and inherent data ambiguities, exacerbates these issues.
  • Algorithmic limitations: Algorithms governing LLM input processing and output generation often lack the sophistication to consistently grasp context or verify factual accuracy, leading to contextually inappropriate or factually incorrect responses.
Manifestations. Hallucinations in LLMs manifest in various forms, from minor inaccuracies to entirely fictitious narratives. Sometimes, these manifest as confident but false assertions, particularly misleading when LLMs are employed in sensitive fields such as medical diagnostics [132], legal advising [217], social content moderation [215], or education [247].
Mitigation Strategies. Numerous research has attempted to mitigate hallucination in LLMs [120]. Most existing mitigation strategies can be categorized into data-centric approaches [208, 90, 3, 238, 140, 347] and model-centric approaches [151, 162, 222, 281]. In the data-centric approaches, several works aim to improve the quality of training data, ensuring it is accurate, diverse, and free of biases. This may involve rigorous data curation and validation processes [169]. Tian et al. introduce the external knowledge graph to mitigate the problem of hallucinations [267]. For the model-centric approaches, many works enhance the model architectures for a better understanding of context, discern factual accuracy, and recognize when the model is venturing into areas of low confidence or outside its training scope [134]. This could involve incorporating mechanisms to check factual accuracy in real time or integrating feedback loops that allow the model to learn from its mistakes. Yao et al. directly edit model parameters to bridge the knowledge gap to mitigate hallucinations [317]. While substantial progress has been made in identifying and categorizing hallucinations [155], the development of robust mechanisms to prevent or correct these errors remains an ongoing area of research. This is crucial for LLMs’ future advancements in various fields.
Discussion. Detecting instances when LLMs are prone to hallucinations is crucial. While the bulk of research on LLM hallucination has centered on the English language, it has been shown that these models are more prone to hallucinations in non-English languages [132]. This disparity underscores a significant gap in our understanding of hallucinations within multilingual contexts and underscores the urgency in developing robust detection and mitigation strategies for hallucinations in diverse linguistic environments. Furthermore, most existing studies have been centered around unimodal hallucinations. However, the emergence of multimodal LLMs, capable of synthesizing and interpreting data across different modalities such as text, images, and audio, poses unique challenges [279, 169, 80, 82, 318, 153]. Overall, addressing hallucination effectively in LLMs requires a comprehensive approach that encompasses multiple languages, modalities, and cultural contexts. Furthermore, transparency regarding operational mechanisms and the inherent limitations of models is vital. Educating users about the potential for hallucinations and the specific contexts in which they are most likely to occur can enable a more critical evaluation of outputs generated by LLMs.
3.1.2 Sycophancy
Large language models may exhibit a tendency to flatter users by reaffirming their misconceptions and stated beliefs, a behavior known as sycophancy [122]. This issue raises significant concerns about the model’s ability to provide objective and unbiased information. Sycophancy in LLMs can lead to the reinforcement of incorrect beliefs, limiting the educational and corrective potential of these systems, and potentially exacerbating echo chambers in digital interactions [251, 148].
Underlying Causes. The propensity for sycophancy can be attributed to several factors:
  • Model size: Research indicates that as model sizes increase, such as reaching scales up to 52 billion parameters, the likelihood of exhibiting sycophantic behaviors also rises [261], potentially due to the increased capacity to model and mirror user preferences.
  • Training method: Reinforcement Learning from Human Feedback (RLHF) can also increase sycophancy [261]. RLHF may inadvertently prioritize agreeableness or affirmation of user beliefs, especially if the feedback loop is dominated by users who favor or reward such responses.
  • Conversational scenario: Sycophancy is particularly evident in scenarios where users challenge the model’s outputs or engage in interactions that require the model to adapt or comply with user assertions. In such cases, the model might lean towards agreeability to maintain a smooth and engaging interaction, leading to a higher occurrence of sycophantic responses.
Discussion. Future research directions to investigate and resolve the issue of sycophancy in LLMs should focus on several key areas. Firstly, developing methods for detecting when an LLM is likely to be reinforcing misconceptions is crucial. This involves enhancing the model’s ability to recognize and differentiate between fact-based assertions and user opinions. Secondly, there is a need to design algorithms that can introduce a balance between user engagement and factual integrity. These algorithms would ensure that while user interactions remain engaging, they do not compromise on delivering accurate and unbiased information. Moreover, exploring the implementation of feedback mechanisms where users can flag responses perceived as overly agreeable or flattering could provide valuable data for training more objective models. Lastly, interdisciplinary research incorporating insights from psychology and ethics could guide the development of LLMs that maintain a neutral stance, particularly in sensitive or polarized topics. These efforts are essential for advancing LLM technology to be both useful and ethically responsible.
3.2 Social Norm
Social norms play a pivotal role in defining acceptable behavior within societies and significantly influence the behavior of large language models (LLMs). Despite their promising capabilities, LLMs can sometimes produce content that is rude, disrespectful, or unreasonable—attributes collectively referred to as “Toxicity” [261, 298]. This issue not only covers the explicit generation of hate speech, insults, profanities, and threats but also includes more subtle forms of harm, such as ingrained or distributional biases. The presence of toxic outputs can have detrimental effects on individuals, specific groups, and the broader societal fabric, posing a multifaceted challenge in both the development and deployment of these AI systems [298]. Such challenges underscore the need for careful consideration of the ethical implications and societal impacts of LLMs in technological advancement. Toxicity mitigation in LLMs involves aligning the models’ outputs with social norms and values, a process essential for minimizing the generation of harmful content [299]. Alignment is one of the fundamental toxicity mitigation approaches, which not only addresses overt expressions of toxicity but also reduces subtler biases [221].
What is alignment in LLMs and why is it needed? With the transformative evolution in Natural Language Processing (NLP) research and development, the impact and success of large language models (LLMs) [338, 336, 57, 329, 266, 5, 273, 274] has been exceptional, exemplified by ChatGPT [303] developed by OpenAI. One key driver for the popularity and usability of recent LLMs is alignment. Alignment is a technique that aims to ensure that generated responses comply with human values. An example is illustrated in Figure 7. Currently, the standard procedure for aligning large language models (LLMs) primarily includes two approaches: SFT (Supervised Fine-Tuning) [221] and RLHF (Reinforcement Learning from Human Feedback) [55, 17]. Since LLMs have been used in a wide range of applications (e.g., editing/writing assistance, personal consultation, question answering, and customer support), many corresponding concerns would arise if the LLMs are not properly aligned otherwise.
Figure 7: An example of unaligned and aligned LLM response. The unaligned LLM response is problematic because it encourages risky financial behavior without considering the potential downsides. It fails to account for the ethical and responsible aspects of financial advice, potentially leading users to make harmful decisions.
The existing literature suggests various considerations for alignment tasks regarding ethical and social risks [296], however, there is a lack of unified discussion. One general guideline stresses that alignment should be Helpful, Honest, and Harmless, known as the “HHH” principle [11]. Furthermore, Liu et al. [178] present a fine-grained taxonomy of concerns related to unaligned LLMs. In this taxonomy, they categorize the existing works into several aspects, such as fairness, reliability, robustness, explainability, safety, etc.
To address the diverse range of concerns associated with alignment tasks, it is essential to gain a comprehensive understanding of the characteristics of LLM alignments and the corresponding evaluation methods. Subsequently, we study and review recent advances in LLM alignments.
Characteristics of Alignment. To understand the characteristics of LLMs, a diverse array of benchmarks have been introduced [165, 284, 282, 283]. In contrast to general-purpose evaluation, alignment-focused evaluation depends on the taxonomy of alignment, associated with corresponding scenarios, criteria, and datasets [286, 350, 349]. Obtaining appropriate criteria and datasets for evaluating alignments in LLMs is crucial, albeit a non-trivial task  [51, 283]. This essentially involves representing the preferences of humans [41]. However, manually collecting human judgment can be expensive, time-consuming, and labor-intensive [345]. To address this issue, researchers proposed to use strong LLMs as an automated proxy for evaluating other LLMs [348]. For example, AUTO-J [154] is trained to tackle challenges in evaluating LLM alignments regarding generality, flexibility, and interoperability. AUTOCALIBRATE presents a multi-stage, gradient-free approach [154], to automatically calibrate and align an LLM-based evaluator toward human preference free of human intervention.
Recent Advancements in Alignment. In the endeavor to align LLMs with human values, a myriad of research initiatives [263, 244, 257, 302, 313, 129, 248] have been undertaken to achieve effective LLM alignments. The forefront of these approaches emphasizes the generative capabilities of large language models (LLMs) for self-regulation with minimal human supervision. SELF-ALIGN [263] proposes a topic-guided, principle-driven approach to autonomously generate responses that are helpful, ethical, and reliable, leveraging the mechanism of in-context learning. Similarly, KNOWNO [244] is a framework for evaluating and aligning the uncertainty in LLM-based planning. Utilizing the theory of conformal prediction, KNOWNO ensures statistical reliability in task completion, thereby minimizing human assistance in complex planning scenarios. Additionally, PRO [257] introduces a response probability ranking method, enhancing the Bradley-Terry comparison model to effectively direct the LLM to favor the most appropriate response. Complementarily, P3O [302] presents a trajectory-wise policy gradient algorithm, which uniquely focuses on comparative rewards instead of traditional reward optimization trained from comparison-based losses.
Discussion. The burgeoning field of LLM alignment, pivotal for the symbiosis of AI and humanity, anticipates transformative discoveries. Emphasizing the importance of AI safety and the seamless integration of AI with human society, prioritizing the alignment of LLMs, with human ethos is essential. As LLMs’ capabilities escalate, the complexity of achieving this alignment intensifies, necessitating increased scientific and technological investment. This demands an exploration of novel strategies in this domain. Foremost, amidst the rapid evolution of LLMs, it is crucial to guarantee their adherence to human ethical standards, which requires more theoretical breakthroughs [301]. In addition, the growing intricacy of AI architectures calls for automated systems capable of assessing and realigning these models [227]. Next, the black-box nature of LLMs also highlights the urgency for clarity and explainability in their alignment processes [343]. Lastly, leveraging adversarial attacks as a method to test and refine the alignment of LLMs emerges as an effective approach for ensuring their conformity to human values [355].
3.3 Law and Regulatory Compliance
Given new-emerging ethical challenges posed by LLMs, there is an increasing demand for effective regulation and oversight of LLMs to ensure their safe and responsible use [45]. Regulation refers to the rules, standards, and principles that govern the development, deployment, and use of LLMs, such as laws, policies, guidelines, or codes of conduct [294, 39, 246]. Oversight refers to the mechanisms, processes, and institutions that monitor, evaluate, and enforce the compliance of LLMs with regulations, such as audits, reviews, certifications, or sanctions [236]. Regulation and oversight of LLMs aim to protect the rights, interests, and values of the stakeholders involved, such as data owners, users, developers, providers, regulators, and society at large [203].
With that being said, the use of LLMs has not yet been resolved by a consensus or a clear regulation therefore posing ethical and legal challenges. European Union (EU) has made substantial efforts in the law and regulations on Artificial Intelligence (AI). In the EU, AI tools, such as LLMs, are subject to the General Data Protection Regulation (GDPR), which regulates the collection, processing, and analysis of personal data, as well as automated decision-making that affects individuals [241]. In this sense, for a company to operate lawfully in the EU regarding the collection and processing of personal data, it must follow the principles and rules laid down in the GDPR. Furthermore, on May 13, 2022, the French Council presidency circulated an amendment to the draft AI Act  [^1] https://data.consilium.europa.eu/doc/document/ST-14954-2022-INIT/en/pdf, on what the text calls “general-purpose AI systems” (GPAIS) [26, 27]. This novel passage has come to form the nucleus of the direct regulation of LLMs and contains rules on the AI value chain [28].
On 30 March 2023, the Italian Data Protection Authority ordered the temporary suspension of the processing of personal data of subjects established on Italian territory by OpenAI LLC, a US company that develops and manages ChatGPT, because the chatbot had failed to comply with the rules set out in GDPR, as well as the Italian Personal Data Protection Code [225]. Meanwhile, the EU parliament is continuously working on the EU AI Act, which is poised to be the World’s first regulation on AI [125]. This Act envisions a distinct regulatory framework compared to the proposals under consideration in the United Kingdom and categorizes AI systems based on varying risk levels, enabling tailored regulations that correspond to each level of risk [275]. At the time of writing this manuscript, several other countries are exploring the possibility of limiting or regulating the use of LLMs [86, 163].
Discussion. Despite the heroic striving of the AI Act to keep up with the accelerating dynamics of AI development, several discussions are also proposed around its practical compliance with LLMs. Hacker et al. argue that this direct regulation is unsatisfactory and could be further enhanced from 1) the definition of GPAIS, 2) the risk management of GPAIS, and 3) the adverse consequences for competition [95]. They propose to focus on the deployers and users more and directly apply non-discrimination and data protection law (GDPR compliance) on LLMs. Bommasani et al. [34] systematically evaluate the compliance with the draft EU AI Act of the foundation model providers like OpenAI and Google. They evaluate the compliance of 10 major foundation model providers (and their flagship models) with the 12 requirements proposed by the EU AI Act and use a scale from 0 (worst) to 4 (best) to rate each provider and model for each requirement. The best possible score for a provider or a model is 48, which indicates full compliance with the AI Act. Their results identify four areas where many organizations receive low scores (usually 0 or 1 out of 4) in terms of compliance with the AI Act: 1) copyrighted data, 2 compute/energy, 3) risk mitigation, and 4) evaluation/testing. Aside from these general regulations, there are also discussions on challenges of how to regulate LLMs for vertical domains such as medical usage [204] and healthcare [203].
4 Conclusion
While presenting remarkable opportunities for advancing artificial intelligence (AI) techniques, Large Language Models (LLMs) expose significant ethical challenges that must be meticulously addressed. Exploring the techniques of LLMs within ethical boundaries is a paramount and complicated endeavor, requiring continual innovation in evolving technological capabilities and societal expectations. In this paper, we survey ethical issues posed by LLMs from longstanding challenges, such as privacy, copyright, and fairness, to new-emerging dilemmas related to truthfulness, social norms, and regulatory compliance. We also discuss the existing approaches that mitigate the potential ethical risks and the corresponding future directions. Our survey is a stepping stone for researchers to advance LLM techniques under ethical standards, ensuring positive contributions to our society.